High Desert Potteryworks makes a variety of pottery products that it sells to retailers. The company uses a job-order costing system in which departmental predetermined overhead rates are used to apply manufacturing overhead cost to jobs. The predetermined overhead rate in the Molding Department is based on machine-hours, and the rate in the Painting Department is based on direct labor-hours. At the beginning of the year, the company provided the following estimates:

Department
Molding Painting
Direct labor-hours 12,000 60,000
Machine-hours 70,000 8,000
Fixed manufacturing overhead cost $ 497,000 $ 615,000
Variable manufacturing overhead per machine-hour $ 1.50 -
Variable manufacturing overhead per direct labor-hour - $ 2.00

Job 205 was started on August 1 and completed on August 10. The company's cost records show the following information concerning the job:

Department
Molding Painting
Direct labor-hours 30 84
Machine-hours 110 20
Direct materials $ 770 $ 1,332
Direct labor cost $ 525 $ 1,470

Required:

1. Compute the predetermined overhead rates used in the Molding Department and the Painting Department.

2. Compute the total overhead cost applied to Job 205.

3-a. What would be the total manufacturing cost recorded for Job 205?

3-b. If the job contained 50 units, what would be the unit product cost?
Answer #1
Calculation of pre determined overhead rate
Molding Painting
Fixed manufacturing overhead cost 497,000 615,000
Direct labor hours     12,000     60,000
Machine hours     70,000       8,000
Fixed manufacturing overhead rate         7.10       10.25
Variable manufacturing overhead rate         1.50         2.00
Pre determined overhead rate         8.60       12.25
2.Overhead cost applied = 110*8.6 + 84*12.25 1975
3.Total manufacturing cost
Direct material       2,102
Direct Labor       1,995
Overhead cost       1,975
Total manufacturing cost       6,072
3-B Unit product cost = 6072/50 = $121.44